OpenStudy (anonymous):

Last month, Edwin had $2325.35 that he divided into two bank accounts. He deposited into a savings account. He deposited the rest into a checking account. How much does edwin Deposit in his checking account ? I will medal you

OpenStudy (anonymous):

We need to find what two fifths of 2325.35 is. So this is fairly straight-forward, we are going to divide it by 5 (hence 5ths) and use that value to find the amount deposited into the checking account. So...\[2325.35/5=465.07\]So we know 1/5th of 2325.35 is 465.07.

OpenStudy (anonymous):

We know that Edwin deposited 2/5ths of that into a savings account, so whatever is left over is what went to the checkings account. So, 2 times 465.07 will get us our savings value:\[465.07*2=930.14\] This is the amount that went into the savings, so now we subtract that from the original amount to find the amount deposited into the checking.\[2325.35-930.14=1395.21\] So he deposited $1395.21 into the checking account.